OpenStudy (anonymous):

How to find the arc length of y=1+6x^(3/2), 0

OpenStudy (anonymous):

S = √ ( 1 + (dy/dx)² dy/dx = (6 * 3/2 )√ x = 9√x -> 1 + (dy/dx)² = 81x + 1 Thus S = √ ( 81x +1

OpenStudy (anonymous):

Thank You!